At the beginning of 2017, Bonita Industries issued 8% bonds with a face value of $4900000. These bonds mature in five years, and interest is paid semiannually on June 30 and December 31. The bonds were sold for $4539360 to yield 10%. Bonita uses a calendar-year reporting period. Using the effective-interest method of amortization, what amount of interest expense should be reported for 2017? (Round your answer to the nearest dollar.)

$454138
$468352
$455484
$452806
Answer #1

Interest paid semiannually = 4900000*4% = 196000

Interest expense on June 30 = 4539360*5% = 226968

Interest expense on Dec 31 = (4539360+30968)*5% = 228516

Total interest expense 2017 = 228516+226968 = $455484

So answer is c) $455484